The parallel plates in a capacitor, with a plate area of 5.30 cm2 and an air-filled separation of 4.60 mm, are charged by a 3.60 V battery. They are then disconnected from the battery and pulled apart (without discharge) to a separation of 6.00 mm. Neglecting fringing, find (a) the potential difference between the plates, (b) the initial stored energy, (c) the final stored energy, and (d) the work required to separate the plates.
Answer #1

Part A.

Capacitance of a capacitor is given by:

C1 = e0*A1/d1

Charge across this capacitor will be

Q1 = C1*V1

Q1 = e0*A1*V1/d1

Now when plates are pulled apart without discharging, then Charge and surface area of plates will remain constant, So

Q2 = e0*A2*V2/d2

Since

Q1 = Q2 & A1 = A2

Q1 = e0*A1*V2/d2

V2 = Q1*d2/(e0*A1)

V2 = d2*Q1/(d1*e0*A1/d1)

V2 = (d2/d1)*Q1/C1 = (d2/d1)*V1

Now d2 = 6.00 mm, d1 = 4.60 mm, & V1 = 3.60 V

So,

V2 = (6.00/4.60)*3.60

V2 = 4.70 V

Part B.

Initial energy stored will be:

U1 = (1/2)*C1*V1^2

U1 = (1/2)*e0*A1*V1^2/d1

Using given values:

U1 = (1/2)*(8.85*10^-12*5.30*10^-4*3.60^2/(4.60*10^-3))

U1 = 6.61*10^-12 J = 6.61 pJ

Part C.

final energy stored will be:

U2 = (1/2)*C2*V2^2

U2 = (1/2)*e0*A2*V2^2/d2

Since A1 = A2

Using given values:

U2 = (1/2)*(8.85*10^-12*5.30*10^-4*4.70^2/(6.00*10^-3))

U2 = 8.63*10^-12 J = 8.63 pJ

Part D.

Work-done to separate these plates will be:

W = Change in potential energy

W = U2 - U1

W = (8.63*10^-12) - (6.61*10^-12)

W = 2.02*10^-12 J

Please Upvote.